OpenStudy (anonymous):

I keep getting 8.2 for my answer, but the book says 5. Could someone explain where I'm going wrong? 1/4 (x+3) - 2/3 (x-1) = -2 Solve for x.

OpenStudy (anonymous):

Some easier to read formatting: \[\frac{ 1 }{ 4 }(x+3) - \frac{ 2 }{ 3 } (x-1) = -2\]

OpenStudy (anonymous):

what do u have some far when u distribute ?

OpenStudy (anonymous):

Well, I started out by multiplying everything by 12 to get rid of the fractions: \[3(x+3)-8(x-1)=-24\]

OpenStudy (anonymous):

Then distributing: \[3x+9-8x+8=-24\] Combining stuff: \[-5x+17=-24\] \[-5x=-41\] =8.2

OpenStudy (anonymous):

can see what u did wrong!!! looks right to me maybe wrong answer in back of book happened to me before

OpenStudy (anonymous):

so lets try plugging it into the equation and see if we get the right answer

OpenStudy (anonymous):

8.2 is correct if put into the original equation

OpenStudy (anonymous):

That's what I'm figuring. It's for a friend, she came to me for help and I couldn't give her the same answer as her book had! Plugging it back in gets me -2

OpenStudy (anonymous):

thats what i love about math can always check your answer back into the equation!!

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }x+\frac{ 3 }{ 4 }-\frac{ 2 }{ 3 }x+\frac{ 2 }{ 3 }=-2\] \[12(\frac{ 1x }{ 4 })+12(\frac{ 3 }{ 4 })-12(\frac{ 2x }{ 3 })+12(\frac{ 2 }{ 3 })=12(-2)\] \[3x+9-8x+8=-24\] \[-5x+17=-24\] \[-5x=-41\] \[x=\frac{-41}{-5}\] \[x=8.2\] So in short, you are definitely doing it right. Maybe you've somehow written down the problem wrong and that's why the answers don't match??

OpenStudy (anonymous):

Book was incorrect.